According to the U.S. Census Bureau, among males over the age of 24, 13.4% did not complete high school, 31.9% completed high school, and 16.5% attended some college without graduating. (Source: U.S. Department of Commerce, Census Bureau, Current Population Survey Data on Educational Attainment.) a. Compute and interpret the probability that a randomly selected male in the United States who is over the age of 24 did not attend college.
Answer #1

Did not complete high school = 13.4%

Completed high school = 31.9%

Attended some college without graduating = 16.5%

Graduation and above = (100 - 13.4+31.9+16.5) = 38.2%

a) P(a randomly selected male in the United States who is over the age of 24 did not attend college) = 0.134 + 0.319

= 0.453

45.3% of males in the United States who is over the age of 24 did not attend college